OpenStudy (liv33):

Three friends go bowling. The cost per person per game is $5.30. The cost to rent shoes is $2.50 per person. Their total cost is $55.20. How many games did they play? Set up the equation and explain your work, please

OpenStudy (greatlife44):

well first what we can do is this: subtract the cost of the shoes from the total cost 1. $2.50*(3) = 7.50 2. $55.20-($7.50) = $47.70

OpenStudy (greatlife44):

\[\frac{ $47.70 }{ 5.30 } = number~of~games~played \]

OpenStudy (liv33):

So, around 8 games?